OpenStudy (anonymous):

A long distance telephone company charges a rate of 6 cents per minute or a 50- cent minimum charge per completed call, whichever is greater. Find the costs, in cents per minute, for a 5 minute call. (ROUND TO A NEAREST TENTH OF A CENT)

OpenStudy (anonymous):

the equation is 6t + 50 = cost with t being time does that help?

OpenStudy (anonymous):

t=minutes yes

OpenStudy (anonymous):

huh still at loss

OpenStudy (anonymous):

80

OpenStudy (anonymous):

hold on no the equation is wrong let me start from the top

OpenStudy (mathstudent55):

If it costs 6 cents per minute, how much does a 5-minute call cost at that rate?

OpenStudy (anonymous):

5(6*60)=cost

OpenStudy (mathstudent55):

@dolloway97 If it costs 6 cents per minute, how much does a 5-minute call cost at that rate?

OpenStudy (anonymous):

oh wow its per minute ive been reading it as per second this whole time

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

30 is the cost

OpenStudy (mathstudent55):

1 minute costs 6 cents 2 minutes cost 2 * 6 cents = 12 cents 3 minutes cost 3 * 6 cents = 18 cents What about 5 minutes?

OpenStudy (anonymous):

5*6=?

OpenStudy (anonymous):

i said 30

OpenStudy (mathstudent55):

Right. At the rate 6 cents per minute, a 5-minute call would cost 30 cents. Now lets; read carefully the problem.

OpenStudy (mathstudent55):

The problem states: "A long distance telephone company charges a rate of 6 cents per minute or a 50-cent minimum charge per completed call, whichever is greater."

OpenStudy (mathstudent55):

From this part: "A long distance telephone company charges a rate of 6 cents per minute" you'd have a 30-cent charge for a 5-minute call, but from this part: "... or a 50-cent minimum charge per completed call, whichever is greater." you see that the company will never charge you less than 50 cents for any call. Since a 5-minute call is only 30 cents, the company charges the 50-cent minimum for it. That means a 5-minute call really costs 50 cents. Do you understand this?

OpenStudy (anonymous):

but its 50 cents minimum charge per completed call

OpenStudy (mathstudent55):

Right. That is why even though this call should have cost 30 cents, it really costs 50 cents. Any call that costs less than 50 cents has its price raised to 50 cents because of the 50-cent minimum charge the company has.

OpenStudy (anonymous):

so since its 50 cents for a five minute call what was the sense in finding out 6x5=30

OpenStudy (anonymous):

but its asking HOW MUCH DOES A 5 MIN CALL COST IN CENTS PER MINUTE

OpenStudy (mathstudent55):

A great question. We needed to see if the charge was going to be the actual cost of the call or the minimum 50-cent charge. We needed to see whether 5 minutes at 6 cents per minute would by more or less than the 50-cent minimum. Once we multiplied 5 * 6 and got 30 cents, now we know we use the minimum charge and not eh 30-cent charge.

OpenStudy (mathstudent55):

We're getting there. Now that we know the charge is 50 cents for a 5 minute call, we finish the problem.

OpenStudy (mathstudent55):

Since the 5-minute call really cost 50 cents, now we can find the cost per minute for the 5-minute call: \(cost ~per ~minute = \dfrac{cost ~of ~the ~call}{number ~of ~minutes}\) \(cost ~per ~minute = \dfrac{50~cents}{5~minutes}\) What is 50/5 = ?

OpenStudy (anonymous):

SO A 5 MINUTE CALL COSTS 50 CENTS PER MINUTE

OpenStudy (anonymous):

50/5=10 @mathstudent55

OpenStudy (mathstudent55):

Right. A 5-minute call costs 50 cents, and since 50 cents/5 minutes = 10, it costs 10 cents per minute.